OpenStudy (baby456):

Could someone Please Help Me! 1. (x-7)^2=81 2. ( x-5)(3x-2)=0

OpenStudy (baby456):

@jdoe0001

OpenStudy (baby456):

@wolf1728

OpenStudy (baby456):

@jim_thompson5910

OpenStudy (baby456):

@jigglypuff314

OpenStudy (baby456):

@zzr0ck3r

OpenStudy (zzr0ck3r):

\((x-7)^2=81 \implies x-7=\pm\sqrt{81}=\pm 9 \implies x=\pm9+7\)

OpenStudy (mathmale):

Regarding solving ( x-5)(3x-2)=0: Set each factor = to 0 separately and solve each of the resulting equations for x. You'll obtain two solutions by doing this.